OpenStudy (anonymous):

Variable y varies directly with x^2, and y = 196 when x = 7. Enter the constant of variation in the quadratic variation equation. y = x2

OpenStudy (dtan5457):

y=kx That is your formula for direct variation. y=x^2 Given that y=196, and x=7. 196=(7^2)(K) Just solve for k, or the constant.

OpenStudy (anonymous):

How would i solve for "K"

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Well, I mean I plugged everything in for you. 196=49(k) What would k be to make the equation true?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

Yes, 4 would be your constant.

OpenStudy (anonymous):

Okay, So how would I put it in?

OpenStudy (dtan5457):

Not sure what your asking. y=kx y and x is given. k is 4.

OpenStudy (anonymous):

y=___x^2 so does 4 go where the ___ is?

OpenStudy (dtan5457):

y=4x^2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Can you help with this? http://brainly.com/question/264009

OpenStudy (dtan5457):

Welcome. Feel free to fan and medal if that helped. Tag me in any more questions if needed.

OpenStudy (dtan5457):

Sure, give me a sec,

OpenStudy (anonymous):

Thank you so so so much!

OpenStudy (dtan5457):

Well this function says that y=x^2 y=18, x=6 18=k(6^2) 18=k(36) What is k?

OpenStudy (dtan5457):

Or, what times 36 equals 18?

OpenStudy (anonymous):

2?

OpenStudy (dtan5457):

Haha, it's 1/2. 36 x 2=72 36 x 1/2=18

OpenStudy (dtan5457):

So now that you have 1/2. Your function is now y=1/2x^2

OpenStudy (anonymous):

OHHH okay!!! So what graph is it?

OpenStudy (anonymous):

b?

OpenStudy (dtan5457):

Well if the function is y=1/2x^2 What do you think? It's easy if you plug in any x value and see if the y value comes out as the graph you think it is..

OpenStudy (anonymous):

B?

OpenStudy (dtan5457):

Yes :)

OpenStudy (anonymous):

Yay!!! thank you!!!

OpenStudy (dtan5457):

No problem. Tag me in anymore questions if you have any.

OpenStudy (anonymous):

Okay thank you:)